# Is Complex Query Answering Really Complex?

### v2

@@ A.3 Influence of intermediate existing entities on the results
For query types having intermediate variables, i.e., 2p, 3p, 2i1p, 1p2i, and 2u1p, we analyzed how the number of existing intermediate entities matching the existentially quantified variables could influence the results. We refer to this number as cardinality of existing entities. For example, in Figure 1 the entities “When in Rome” and “Spiderman 2” are existing intermediate entities matching the existential quantified variable $V$ for the query $q_{1}$. The intuition is that while some of these entities simplify the prediction of the answers of the query (see Sec. 3), some others might be misleading for the prediction. Hence, if there is a high number of existing entities in the path, it would be harder for the model to select only the existing entities needed to predict the next link. For this reason, we claim that a high value of cardinality of existing entities would make a $(q,a)$ pair harder.
To support our claim, we analyze the percentage of $(q,t)$ pairs having different values of cardinality, shown in Figure A.3 , and their MRR, using CQD (Arakelyan et al., 2021), shown in Figure A.3. Figure A.3 shows that CQD results are highly influenced by the value of cardinality, having decreasing performance at the increase of the cardinality. Note that by matching this plot with the one in Figure A.3, one can understand if results on a specific query type are highly influenced by a high proportion of $(q,t)$ pairs having the same cardinality. For example, the MRR of 1p2i partial-inference $(q,t)$ pairs in $\mathsf{FB15k237}$ is highly influenced by the presence of about 30% of $(q,t)$ pairs having cardinality of “1”.
@@
Figure A.3: Proportion of$(q,t)$ pairs having different cardinality of intermediate existing entities in the old benchmarks.
@@ D.1 Influence of intermediate existing entities on the results
Similarly to what was done for the old benchmarks in App. A, we analyze the proportion and influence (on the MRR) of the number of intermediate existing entities for CQD, on the new benchmarks $\mathsf{FB15k237}$+$\mathsf{H}$, $\mathsf{NELL995}$+$\mathsf{H}$, $\mathsf{ICEWS18}$+$\mathsf{H}$. Figure D.2 show a very similar result to the one in Figure A.3, showing a decreasing MRR at the increase of the value of cardinality for all benchmarks. Moreover, for reference, we also report the proportion of the different $(q,t)$ pairs for each cardinality category in Figure D.2.
@@
Figure D.2: Proportion of$(q,t)$ pairs having different cardinality of intermediate existing entities in new benchmarks.
@@
When creating the new benchmarks $\mathsf{FB15k237}$+$\mathsf{H}$, $\mathsf{NELL995}$+$\mathsf{H}$, $\mathsf{ICEWS18}$+$\mathsf{H}$, to make sure that no anchor entities nor relation name was predominant in each query type, we set a maximum of 20% of $(q,a)$ pairs having the same anchor entity or relation name across all benchmarks and query type, as shown in Table D.1
@@
For reference, we also provide the stratified analysis on the new benchmarks in Tables A.6 and A.5, which shows that 1) the full-inference QA pairs have always a lower MRR than the overall, and 2) the overall MRR is much lower than the 1p-reductions only, as for the new benchmark the proportion of QA pairs is equally distributed.
@@ Appendix F Hyperparameters
In this section, we detail the hyperparameters used for each dataset and model. Old and new benchmarks, the generation scripts, and the implementation of CQD-hybrid are included as supplemental materials in a GitHub repo. [^6] https://github.com/april-tools/is-cqa-complex/tree/main
@@ F.1 Old benchmarks
As mentioned in Sec. 5, we re-used the pre-trained link predictor (Trouillon et al., 2017) provided by the authors. [^9] https://github.com/Blidge/KGReasoning/ However, we tuned CQD-specific hyperparameters, namely the CQD beam “k”, ranging from [2,512] and the t-norm type being “prod” or “min” In Table F.1 we provide the hyperparameter selection for the old benchmarks $\mathsf{FB15k237}$ and $\mathsf{NELL995}$. Also note that we normalize scores with min-max normalization.
Dataset 2p 3p 2i 3i 1p2i 2i1p 2u 2u1p
k tn k tn k tn k tn k tn k tn k tn k tn
$\mathsf{FB15k237}$ 512 p 8 p 128 p 128 p 256 p 64 p 512 m 512 m
$\mathsf{NELL995}$ 512 p 2 p 2 p 2 p 256 p 256 p 512 m 512 m
@@ F.2 New benchmarks
For $\mathsf{FB15k237}$+$\mathsf{H}$ and $\mathsf{NELL995}$+$\mathsf{H}$, we re-used the same models trained for the old benchmarks, and using the same hyperparameters presented in Sec. F.1. Instead, for the new benchmark $\mathsf{ICEWS18}$+$\mathsf{H}$ we trained every model. The used hyperparameters for each model are presented in the following:
@@
For $\mathsf{NELL995}$+$\mathsf{H}$ and $\mathsf{FB15k237}$+$\mathsf{H}$ we re-used the pre-trained link-predictor and the same hyperparameters found for the old benchmarks F.1. Instead, for the newly created $\mathsf{ICEWS18}$+$\mathsf{H}$, we train ComplEx (Trouillon et al., 2017) link predictor with hyperparameters “regweight” 0.1 or 0.01, and batch size 1000 or 2000, with the best being, respectively 0.1 and 1000. Moreover, in F.2 are shown the hyperparameters for CQD on the new benchmark $\mathsf{ICEWS18}$+$\mathsf{H}$.
Dataset 2p 3p 2i 3i 1p2i 2i1p 2u 2u1p 4i 4p
k tn k tn k tn k tn k tn k tn k tn k tn k tn k tn
$\mathsf{ICEWS18}$+$\mathsf{H}$ 32 p 2 p 2 p 2 p 512 p 256 p 2 m 2 m 2 p 2 p
Table F.2: CQD hyperparameters new $\mathsf{ICEWS18}$+$\mathsf{H}$ benchmark. tn = tnorm. p=prod, m=min
@@ G.2.1 $\mathsf{NELL995}$+$\mathsf{H}$ and $\mathsf{FB15k237}$+$\mathsf{H}$
queries are generated to ensure a balanced distribution of hardness across the sub-query types to which a given query can be reduced. As described in Sec. 6, we generate queries such that each sub-query type contains exactly 10,000 $(q,t)$ pairs, along with an additional 10,000 pairs corresponding to full-inference queries (i.e., non-reducible cases). To achieve this, we adapt the algorithm from (Ren et al., 2020), sampling queries until the required number of $(q,t)$ pairs for each sub-query type is reached. For instance, for 2p queries, we continue sampling until we obtain 10,000 $(q,t)$ pairs reducible to 1p and 10,000 full-inference queries. When a query is sampled, all its corresponding answers are included in the benchmark until the target of 10,000 $(q,t)$ pairs per subtype is met. If incorporating all answers would exceed this limit, a random subset is selected to ensure exactly 10,000 pairs. Similarly, once a sub-query type reaches the 10,000-pair threshold, any additional answers belonging to that category are discarded.
@@ G.2.2 ICEWS18+H
were generated using the same strategy detailed in Sec. G.2.1. Moreover, since we generated this benchmark from scratch, we also had to generate the 1p, 2in, 2nu1p queries, which following (Ren & Leskovec, 2020; Ren et al., 2020), were generated for each triple in the validation and test split.

### v3

@@ A.3 Influence of intermediate existing entities on the results
For query types having intermediate variables, i.e., 2p, 3p, 2i1p, 1p2i, and 2u1p, we analyzed how the number of existing intermediate entities matching the existentially quantified variables could influence the results. We refer to this number as cardinality of existing entities. For example, in Figure 1 the entities “When in Rome” and “Spiderman 2” are existing intermediate entities matching the existential quantified variable $V$ for the query $q_{1}$. The intuition is that while some of these entities simplify the prediction of the answers of the query (see Sec. 3), some others might be misleading for the prediction. Hence, if there is a high number of existing entities in the path, it would be harder for the model to select only the existing entities needed to predict the next link. For this reason, we claim that a high value of cardinality of existing entities would make a QA pair harder.
To support our claim, we analyze the percentage of QA pairs having different values of cardinality, shown in Figure A.2 , and their MRR, using CQD (Arakelyan et al., 2021), shown in Figure A.2. Figure A.2 shows that CQD results are highly influenced by the value of cardinality, having decreasing performance at the increase of the cardinality. Note that by matching this plot with the one in Figure A.2, one can understand if results on a specific query type are highly influenced by a high proportion of QA pairs having the same cardinality. For example, the MRR of 1p2i partial-inference QA pairs in $\mathsf{FB15k237}$ is highly influenced by the presence of about 30% of QA pairs having cardinality of “1”.
@@
Figure A.2: Proportion ofQA pairs having different cardinality of intermediate existing entities in the old benchmarks.
@@ D.1 Influence of intermediate existing entities on the results
Similarly to what was done for the old benchmarks in App. A, we analyze the proportion and influence (on the MRR) of the number of intermediate existing entities for CQD, on the new benchmarks $\mathsf{FB15k237}$+$\mathsf{H}$, $\mathsf{NELL995}$+$\mathsf{H}$, $\mathsf{ICEWS18}$+$\mathsf{H}$. Figure D.2 show a very similar result to the one in Figure A.2, showing a decreasing MRR at the increase of the value of cardinality for all benchmarks. Moreover, for reference, we also report the proportion of the different QA pairs for each cardinality category in Figure D.2.
@@
Figure D.2: Proportion ofQA pairs having different cardinality of intermediate existing entities in new benchmarks.
@@
When creating the new benchmarks $\mathsf{FB15k237}$+$\mathsf{H}$, $\mathsf{NELL995}$+$\mathsf{H}$, $\mathsf{ICEWS18}$+$\mathsf{H}$, to make sure that no anchor entities nor relation name was predominant in each query type, we set a maximum of 20% of QA pairs having the same anchor entity or relation name across all benchmarks and query type, as shown in Table D.1
@@
For reference, we also provide the stratified analysis on the new benchmarks in Tables A.5 and A.4, which shows that 1) the full-inference QA pairs have always a lower MRR than the overall, and 2) the overall MRR is much lower than the 1p-reductions only, as for the new benchmark the proportion of QA pairs is equally distributed.
@@
Model CLMPT CQD CQD-h ConE GNNQE QTO ULTRA
CLMPT 3.32e-63 1.82e-80 3.92e-01 8.28e-71 1.06e-188 8.55e-35
CQD 3.32e-63 1.48e-04 1.94e-52 1.72e-01 5.27e-38 2.31e-06
CQD-h 1.82e-80 1.48e-04 8.88e-75 5.08e-03 8.41e-15 6.91e-16
ConE 3.92e-01 1.94e-52 8.88e-75 6.65e-60 1.38e-165 7.04e-28
GNNQE 8.28e-71 1.72e-01 5.08e-03 6.65e-60 4.79e-30 2.32e-09
QTO 1.06e-188 5.27e-38 8.41e-15 1.38e-165 4.79e-30 2.22e-70
ULTRA 8.55e-35 2.31e-06 6.91e-16 7.04e-28 2.32e-09 2.22e-70
Table D.2: Pairwise Mann–Whitney U test $p$-values on reciprocal ranks of 2p queries in $\mathsf{FB15k237}$+$\mathsf{H}$. Lower values indicate stronger evidence of a significant difference between model performances.
@@
CLMPT 2.5 3.4 2.4 4.5 6.3 3.8 0.2 0.1 0.2
D.2 Statistical tests
As an illustrative case, we analyze 2p queries from $\mathsf{FB15k237}$+$\mathsf{H}$ to assess whether the observed differences in model performance are statistically significant, despite similar MRR in Table 5. We conduct an all-vs-all paired Mann–Whitney U test on the reciprocal ranks across all models for each 2p QA pair. The results in Table D.2 reveal strong statistical differences in most pairwise comparisons, indicating that models rank correct answers differently even when aggregate metrics are close. However, a few model comparisons—such as ConE vs. CLMPT and CQD vs. GNNQE—exhibit statistical similarity, suggesting comparable ranking behavior on 2p queries. To examine whether the similarity is limited to specific query types, we apply the same test across all QA pairs of $\mathsf{FB15k237}$+$\mathsf{H}$ from all query types. The resulting p-values are effectively zero for all comparisons, confirming that significant performance differences are widespread, though limited cases of similarity may occur between some models on some specific query types.
@@ Appendix F Hyperparameters
In this section, we detail the hyperparameters used for each dataset and model. Old and new benchmarks, the generation scripts, and the implementation of CQD-hybrid are included in our official repo. [^6] https://github.com/april-tools/is-cqa-complex
@@ F.1 Old benchmarks
As mentioned in Sec. 5, we re-used the pre-trained link predictor (Trouillon et al., 2017) provided by the authors. [^9] https://github.com/Blidge/KGReasoning/ However, we tuned CQD-specific hyperparameters, namely the CQD beam “k”, ranging from [2,512] and the t-norm type being “prod” or “min” In Table F.1 we provide the hyperparameter selection for the old benchmarks $\mathsf{FB15k237}$ and $\mathsf{NELL995}$. Also note that we normalize scores with min-max normalization. For negation, we always use standard negation.
QTO
We re-used the hyperparameters provided in the official repo.  [^10] https://github.com/bys0318/QTO/tree/main For a fair comparison with CQD, CQD-hybrid, we used the same pre-trained link predictor.
CLMPT
We did not tune hyperparameters but re-used the ones provided in the official repo. [^11] https://github.com/qianlima-lab/clmpt
Dataset 2p 3p 2i 3i 1p2i 2i1p 2u 2u1p 2in 3in 2pi1pn 2nu1p 2in1p
k tn k tn k tn k tn k tn k tn k tn k tn k tn k tn k tn k tn k tn
$\mathsf{FB15k237}$ 512 p 8 p 128 p 128 p 256 p 64 p 512 m 512 m 2 p 2 p 512 prod 512 prod 512 prod
$\mathsf{NELL995}$ 512 p 2 p 2 p 2 p 256 p 256 p 512 m 512 m 2 p 2 p 512 prod 512 prod 512 prod
@@ F.2 New benchmarks
For $\mathsf{FB15k237}$+$\mathsf{H}$ and $\mathsf{NELL995}$+$\mathsf{H}$, we re-used the same models trained for the old benchmarks, and using the same hyperparameters presented in Sec. F.1. The only exception is the hyperparameters for CQD and CQD-hybrid of 4p and 4i queries that were not tuned for the old benchmarks. We used k=2 and tnorm=prod for both query types in all benchmarks.
Instead, for the new benchmark $\mathsf{ICEWS18}$+$\mathsf{H}$ we trained every model. The used hyperparameters for each model are presented in the following:
@@
We train ComplEx (Trouillon et al., 2017) link predictor with hyperparameters “regweight” 0.1 or 0.01, and batch size 1000 or 2000, with the best being, respectively 0.1 and 1000. Moreover, in F.2 are shown the hyperparameters for CQD on the new benchmark $\mathsf{ICEWS18}$+$\mathsf{H}$.
QTO
We re-used the link predictor used in CQD and CQD-Hybrid, and for QTO-specific hyperparameters, we used the same of $\mathsf{FB15k237}$.
Dataset 2p 3p 2i 3i 1p2i 2i1p 2u 2u1p 4i 4p 2in 3in 2pi1pn 2nu1p 2in1p
k tn k tn k tn k tn k tn k tn k tn k tn k tn k tn k tn k tn k tn k tn k tn
$\mathsf{ICEWS18}$+$\mathsf{H}$ 32 p 2 p 2 p 2 p 512 p 256 p 2 m 2 m 2 p 2 p 2 p 2 p 512 prod 512 prod 512 prod
Table F.2: CQD hyperparameters $\mathsf{ICEWS18}$+$\mathsf{H}$ benchmark. tn = tnorm. p=prod, m=min
CQD-Hybrid
For CQD-Hybrid, as mentioned above, we re-used the hyperparameters used for CQD. Note that for 4p queries, we fixed the “k” upper bound to 64 for memory constraints.
@@
CLMPT
For CLMPT, we tuned the following hyperparameters: (1) learning rate, with values in [1e-5,5e-2,5e-3,5e-4,5e-5,,5e-6], (2) temp, with values in [0.1, 0.2] We found the best hyperparameters to match the same found for $\mathsf{FB15k237}$. Hence, we re-used them for $\mathsf{ICEWS18}$+$\mathsf{H}$.
@@ G.2.1 $\mathsf{NELL995}$+$\mathsf{H}$ and $\mathsf{FB15k237}$+$\mathsf{H}$
queries are generated to ensure a balanced distribution of hardness across the sub-query types to which a given query can be reduced. As described in Sec. 6, we generate queries such that each sub-query type contains exactly 10,000 QA pairs, along with an additional 10,000 pairs corresponding to full-inference queries (i.e., non-reducible cases). To achieve this, we adapt the algorithm from (Ren et al., 2020), sampling queries until the required number of QA pairs for each sub-query type is reached. For instance, for 2p queries, we continue sampling until we obtain 10,000 QA pairs reducible to 1p and 10,000 full-inference queries. When a query is sampled, all its corresponding answers are included in the benchmark until the target of 10,000 QA pairs per subtype is met. If incorporating all answers would exceed this limit, a random subset is selected to ensure exactly 10,000 pairs. Similarly, once a sub-query type reaches the 10,000-pair threshold, any additional answers belonging to that category are discarded.
@@ G.2.2 ICEWS18+H
were generated using the same strategy detailed in Sec. G.2.1. Moreover, since we generated this benchmark from scratch, we also had to generate 1p (which following (Ren & Leskovec, 2020; Ren et al., 2020), were generated for each triple in the validation and test split), and 2in, 2nu1p queries.